y-6<=0 3x-2y<=0 5x+3y>=0 how to graph
Answer by Abdul Hannan(9)

Such linear inequalities are often solved graphically in a good approach
consider linear equations for each inequality
for example
write x+y = 4 instead of x+y =< 4
then draw the straight line represented by the equation x/4 + y/4 = 1
now put the values of the origin in the inequality (not in the equation)
ie (x,y) = (0,0)
x+y =< 4
0+0 =< 4
0 =< 4
which is true , thence the graph of the inequality x+y =< 4 is the region bounded by the straight line x+y = 4 and directed towards the origin
had it been false ,the directed away from the origin
in the same manner , draw the graphs of the rest and the req'd solution is the common intersection of all the graphs